Assume h = 0.7 for a mildly deleterious mutation when homozygous (s = 0.05). In this case, the mutation rate to the deleterious form is 10−4. Assuming random mating, what is the equilibrium frequency for the mutant allele? A) 0.00286, B) 0.000314, C) 0.00116, D) 0.00222, E) 0.00346


Please explain, don't just give the answer.
Answer #1

We are given the following:

Hertiability (h) = 0.7

Relative fitness of deleterious mutation (s) = 0.05

Mutation rate (u) = 0.0001

Equilibrium frequency for the mutant allele (q) = u/hs

= 0.0001/(0.7 X 0.05)

= 0.0001/0.035

= 0.00285

Hence, option (a) is correct.